OpenStudy (anonymous):

you have a bag of nickels and quarters that total $3.15. You have four times as many quarters as you have nickels. How many nickels and quarters do you have?

OpenStudy (decarr432):

Hello and welcome to open study!

OpenStudy (decarr432):

Okay this is pretty simple my friend

OpenStudy (decarr432):

how many quarters in a dollar?

OpenStudy (decarr432):

okay how many in 3 dollars?

OpenStudy (decarr432):

okay we have 12 now how many nickles in 15 cents

OpenStudy (decarr432):

Now whats 12 divided by 3?

OpenStudy (decarr432):

Think you just found your answer freind

OpenStudy (decarr432):

No problem

OpenStudy (decarr432):

Please close the question when you are ready

OpenStudy (mathstudent55):

If you are learning systems of equations, and this problem is meant to be solved with a system of equations, then you can do this. Let the number of nickels = n Let the number of quarters = q A nickel is worth $0.05, and a quarter is worth $0.15 n nickels are worth 0.05n, and q quarters are worth 0.25q All the coins together are worth: 0.05n + 0.25q. We are also told they are worth $3.15, so that gives us our first equation: 0.05n + 0.25q = 3.15 For the second equation, we deal with the number of coins. "You have four times as many quarters as you have nickels." Since you have n nickels, you must have 4n quarters. That gives us our second equation: q = 4n Now we have a system of equations in two unknowns that can be solved to find the numbers of nickels and quarters. 0.05n + 0.25q = 3.15 q = 4n Since the second equation is already solved for q, let's use the substitution method and substitute 4n for q in the first equation: 0.05n + 0.25(4n) = 3.15 0.05n + n = 3.15 1.05n = 3.15 n = 3 q = 4n = 4(3) = 12 Answer: there are 3 nickels and 12 quarters.